# PINNfluence: Influence Functions for Physics-Informed Neural Networks

### v1

@@ Abstract
Recently, physics-informed neural networks (PINNs) have emerged as a flexible and promising application of deep learning to partial differential equations in the physical sciences. While offering strong performance and competitive inference speeds on forward and inverse problems, their black-box nature limits interpretability, particularly regarding alignment with expected physical behavior. In the present work, we explore the application of influence functions (IFs) to validate and debug PINNs post-hoc. Specifically, we apply variations of IF-based indicators to gauge the influence of different types of collocation points on the prediction of PINNs applied to a 2D Navier-Stokes fluid flow problem. Our results demonstrate how IFs can be adapted to PINNs to reveal the potential for further studies.
@@ 1 Introduction
Time and time again, deep learning approaches have proven themselves to be exceptionally capable at solving problems that were considered hard and time-consuming, yielding fast inference times bundled with strong performance. Physics-informed neural networks (PINNs) [23] represent another recent iteration of that trend applied to the realm of partial differential equations (PDEs). By incorporating prior physical knowledge in the form of differential equations [6; 9; 10; 23], they admit a wide range of possible applications such as fluid mechanics, electromagnetics, disease modelling and optics, inter alia [1; 2; 4; 19]. Despite the power of PINNs, there remains the challenge of understanding and improving model behavior when things go wrong [15; 25; 26]. PINNs minimize complex composite losses, which complicates the tracing of poor performance to specific training points or conditions; this is the main motivation for this work.
Influence functions [12] (IFs) systematically assess the contribution of individual training points to a model’s behavior. Recent studies have shown that incorporating physical concepts like temporal causality improves PINN performance [7; 28], highlighting the role of underlying physical principles. By identifying key points, such as those near boundaries or critical regions, we can assess whether the PINN’s learned parameters align with the physics of the problem. For many PDEs, practitioners possess an intuitive understanding of the underlying physics. For example, in magnetostatics, the magnetic field is fundamentally generated by source terms. A well-trained PINN could be expected to reflect this relation. We aim to show how domain experts can use this approach to validate the model by identifying key training points, boundary conditions or physical principles that have the greatest impact on the model’s behavior. This also allows them to estimate the effect of inductive biases and observed data on the model’s predictions. Notably, PINNs also provide a powerful application domain for IFs and other data attribution methods, as models and datasets are usually much smaller than those used in computer vision or natural language processing. This makes calculation of IFs computationally feasible, and the effect of individual data points on the fit of the model is expected to be greater.
@@
(i) We adapt influence functions to PINNs as a tool for improving their interpretability. (ii) In addition, we derive heuristically motivated indicators based on influence functions to showcase how they can be leveraged by domain-knowledgeable users to test and validate PINNs. (iii) We empirically demonstrate their usefulness in disambiguating between three different versions of PINNs applied to a Navier-Stokes problem.
@@ 2 Theoretical Background
introduced in [23], constitute a machine-learning-based approach to solving partial differential equations (PDEs), which are ubiquitous in physical sciences. Let $\Omega\subseteq\mathbb{R}^{n}$ be an open domain and consider the initial boundary value problem (IBVP)
@@
The question being addressed by influence functions [5; 8; 12] is how would the model’s behavior change if certain training points were not present in the training set? By linearly approximating the effect of leave-one-out retraining, IFs provide a possibility of studying a model through the lens of its training data. In order to adapt the influence functions to PINNs we will slightly generalize the approach of [12; 13]. We provide a full derivation and proof, including details on all necessary assumptions, in Section A.1. Suppose we have trained a model by minimizing a loss $\frac{1}{N}\Sigma_{i}L(\boldsymbol{x}_{i};\theta)$ on a training data set $\mathcal{X}=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\}$ and derived an optimal parameter $\hat{\theta}\in\Theta$. Furthermore, consider a function $f:\Omega\times\Theta\to\mathbb{R}$, which depends on the parameter $\theta\in\Theta$ and some other parameter $\boldsymbol{z}\in\Omega$ (such as in [12], the loss $L$ for a single test point $\boldsymbol{z}$). We now aim to approximate $f(\boldsymbol{z};\hat{\theta}^{+})-f(\boldsymbol{z};\hat{\theta})$ if $\hat{\theta}^{+}$ were optimized over an amended training data set that includes an additional training point $\boldsymbol{x}^{+}$. This can be approximated by
$\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}^{+}):=\nabla_{\hat{\theta}}f(\boldsymbol{z};\hat{\theta})\cdot H_{\hat{\theta}}^{-1}\cdot\nabla_{\hat{\theta}}L(\boldsymbol{x}^{+};\hat{\theta})\quad\text{,}$ (3)
where $H_{\hat{\theta}}=\nabla^{2}_{\hat{\theta}}\frac{1}{N}\sum_{i=1}^{N}L(\boldsymbol{x}_{i};\theta)$ denotes the Hessian. This approach can be extended to the addition and removal of multiple training data points: Let $\mathcal{X}^{+}=\{\boldsymbol{x}_{1}^{+},\ldots,\boldsymbol{x}_{m^{+}}^{+}\}$ be the set of training points we want to add and $\mathcal{X}^{-}=\{\boldsymbol{x}_{1}^{-},\ldots,\boldsymbol{x}_{m^{-}}^{-}\}\subseteq\mathcal{X}$ the set we want to remove from our training set. The effect of retraining the model with added and removed training points can then be approximated by ${\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})\coloneqq\sum_{i=1}^{m^{+}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{+})-\sum_{i=1}^{m^{-}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{-})}$. Hence, we can study the influence of single sampling points as well as regions on loss functions, any type of predictions or derived quantities. By studying the influence resulting therefrom, one can compare the model’s behavior with physical intuition. Using this method, we can also estimate the effect of adding and removing entire loss terms for composite losses (see Corollary 1 in the appendix).
@@ 3 Experiments
where $\boldsymbol{u}(x,y)\in\mathbb{R}^{2}$ denotes the velocity field, $p(x,y)\in\mathbb{R}$ represents the pressure, $\rho$ is the density, and $\nu$ is the kinematic viscosity. The domain is given by $\Omega=(x_{\min},x_{\max})\times(y_{\min},y_{\max})\setminus C_{r}$, with $C_{r}$ denoting the cylinder with radius $r$. The boundary conditions at the bottom, top, and cylinder are given by $\boldsymbol{u}=0$. The inflow condition on the left has a parabolic profile.
@@
In this study, we aim to answer the following question using influence functions: Can we identify or test whether the PINN has learned certain aspects of the underlying physical principles of the Navier-Stokes problem? Since predictions alone do not reveal the influence of training data or boundary conditions, we devise a set of heuristic indicators.
We propose two IF-based indicators that are designed to capture aspects of the physical processes inherent to the PDE (4) and to serve as exemplary metrics showcasing how expert users can design tests to validate PINNs against their prior knowledge of the problem. As a first indication metric, we concentrate on the flow direction of the Navier-Stokes problem: the inflow is given at the left boundary of the domain and the outflow lies on the juxtaposed side. Figure 2 shows the influence pattern of a training point close to the cylinder.
@@
As a second indicator, we are interested in the fraction of influence that is associated with a given object in the input domain. In the present Navier-Stokes problem, the cylinder plays a central role in the flow field: it obstructs the flow. Thus, the overall velocity and pressure fields should be heavily influenced by its position. To this end, we devise an object or region identifier, which is designed to capture the influence associated with a given object in the domain (in our case around the cylinder). For the object or region indicator ($\operatorname{OI}$), which sums up the relevance of removing a region $\Xi\subseteq\overline{\Omega}$, we write:
$\operatorname{OI}(\Xi):=\frac{\sum_{\boldsymbol{x}_{\text{train}}\in\Xi}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}{\sum_{\boldsymbol{x}_{\text{train}}}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}\in[0,1]$
@@
A higher value of $\operatorname{OI}$ means that the region in question, such as the cylinder in our case, has an increased influence on the prediction for all other points. This is to be expected for a model that aims to accurately reflect the physics. In this investigation, we choose $\Xi=C_{1.5r}$, meaning that we remove the area with a 1.5 times the radius of the cylinder, see Fig. 3.
@@ Results
Table 1: Mean values for the Direction $\overline{\operatorname{DI}}$ and Region $\overline{\operatorname{OI}}$ Indicators, aggregated over all $\boldsymbol{x_{\text{test}}}$. A higher score indicates a closer alignment with the indicator’s assumptions.
$\overline{\operatorname{DI}}$ $\overline{\operatorname{OI}}(C_{1.5r})$
$f(x,\theta)$ $\phi_{\text{good}}$ $\phi_{\text{bad}}$ $\phi_{\text{broken}}$ $\phi_{\text{good}}$ $\phi_{\text{bad}}$ $\phi_{\text{broken}}$
@@
We evaluate the three different iterations of PINNs on the Navier-Stokes task. Figures 3(a) to 3(c) show heatmaps of how influential individual training points were to the whole test set. For $\phi_{\text{good}}$, the most influential points are distributed close to the inflow, on the outflow and especially around the cylinder. This is expected, as these areas determine the fluid flow. Furthermore, we observe that for $\phi_{\text{bad}}$, the upper left boundary is overly influential. And finally, $\phi_{\text{broken}}$’s influences are again distributed well around the cylinder. In Table 1, the aggregated mean values of the indicators are presented. Both $\phi_{\text{good}}$ and $\phi_{\text{bad}}$ show stronger $\operatorname{DI}$ performance. This suggests that both models have successfully learned to capture the dependence of the fluid flow on preceding points. When it comes to the $\operatorname{OI}$ indicator, this trend is even clearer. Here, the good model achieves superior performance, which suggests that the cylinder here has the biggest absolute influence, reflecting its importance for the fluid flow. It is important to note that these indicators are tools for investigating model behavior. Performing well on an indicator does not imply that the model is entirely correct, but rather that it does not fail in the specific aspect the indicator evaluates. Expert knowledge is essential for designing meaningful indicators, and further investigation is needed to validate their robustness. Additionally, if the indicator’s assumptions are flawed, the results may be misleading and fail to provide sensible insights.
@@ 4 Conclusion
We showcased how influence functions can help to evaluate PINNs. In particular, we formulated heuristically motivated indicators designed to test certain physical concepts. By applying this method to a Navier-Stokes problem, we observed that the best-performing model consistently aligned with these indicators, confirming their utility in assessing model performance. Still, further exploration of IF-based markers is indicated to rigorously validate the broad applicability of the presented approach. Using influence functions to improve PINN training through methods like resampling training points [16], enforcing temporal causality [28], or dynamically reweighing loss terms [3] – combined with IF-based training time techniques like TracIN [22] – presents a promising direction for future research. Additionally, further studies on the influence on parts of the loss as well as on different physical quantities, e.g. the vorticity, appears to be a logical next step.
@@ A.1 Influence Functions
In this part, we aim to formulate the key steps for influence functions [8] and adjust them for the application to PINNs. Therefore, we will slightly generalize the results from [12; 13]. Thereby, we will provide proofs for the statements from Section 2. We start with a well-known result [5], which is the crucial ingredient for influence functions.
Lemma 1 (’argmin’-trick).
Let $g:\Theta\times U\to\mathbb{R}$, with $\Theta\subset\mathbb{R}^{P}$ convex, compact and $U\subseteq\mathbb{R}$, be a function, which is twice continuously differentiable w.r.t. $\theta\in\Theta$ and continuously differentiable w.r.t. $\epsilon\in U$. Furthermore, we assume that $g$ has a positive definite Hessian matrix [^1] Note, that any function with a positive definite Hessian is strictly convex, but the converse does not necessary hold. $H_{\theta}\coloneq\nabla^{2}_{\theta}g(\theta,\epsilon)$ for all $\epsilon\in U$. Then $h(\epsilon)\coloneq\operatorname*{arg\,min}_{\theta\in\Theta}g(\theta,\epsilon)$ is unique, and its derivative is given by
$\frac{\partial h(\epsilon)}{\partial\epsilon}=-H_{\hat{\theta}}^{-1}\cdot\frac{\partial}{\partial\epsilon}\nabla_{\hat{\theta}}g(\hat{\theta},\epsilon)\quad\text{.}$ (7)
where $\hat{\theta}\coloneq h(\epsilon)$ denotes the minimum.
@@
Since the Hessian is positive definite, $g(\theta,\epsilon)$ is strictly convex w.r.t. $\theta$ for any $\epsilon\in U$. Further, any continuous function on a compact domain has a minimum (extreme value theorem). Due to the strict convexity, this minimum will be unique $\hat{\theta}=h(\epsilon)$ for any given $\epsilon$. At the minimum we will have
@@
$\nabla_{\hat{\theta}}g(\hat{\theta},\epsilon)=0\quad\text{.}$ (8)
A derivative of Eq. 8 w.r.t. $\epsilon$ leads to
$0=\nabla^{2}_{\hat{\theta}}g(\hat{\theta},\epsilon)\cdot\frac{\partial h(\epsilon)}{\partial\epsilon}+\frac{\partial}{\partial\epsilon}\nabla_{\hat{\theta}}g(\hat{\theta},\epsilon)\quad\text{.}$ (9)
Since the Hessian is positive definite, we can invert the Hessian and a simple rearrangement of Eq. 9 shows the assertion. ∎
By means of this result, we can approximate how an arbitrary function, depending on the optimal parameters $\hat{\theta}$, changes when adding and removing points to our training set.
Proposition 1.
Consider a total loss function $\mathcal{L}(\mathcal{X};\theta)=\frac{1}{N}\sum_{i=1}^{N}L(\boldsymbol{x}_{i};\theta)$ with training points $\mathcal{X}=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\}$, and denote by $\mathcal{X}^{-}=\{\boldsymbol{x}_{1}^{-},\ldots,\boldsymbol{x}_{m^{-}}^{-}\}\subseteq\mathcal{X}$ training points we want to remove and by $\mathcal{X}^{+}=\{\boldsymbol{x}_{1}^{+},\ldots,\boldsymbol{x}_{m^{+}}^{+}\}$ training points we want to add. We will assume that $\mathcal{L}$ has a positive definite Hessian matrix $H_{\hat{\theta}}=\nabla^{2}_{\hat{\theta}}\mathcal{L}(\mathcal{X};\theta)$. Furthermore, let $f:\Omega\times\Theta\to\mathbb{R}$ be an arbitrary function, depending on the parameters $\theta\in\Theta$ and being twice continuously differentiable w.r.t. $\theta\in\Theta$. Then the effect of adding and removing training points on the function $f$ can be approximated by
$f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}_{\mathcal{X}^{+},\mathcal{X}^{-}})=\frac{1}{N}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})+\mathcal{O}(N^{-2})$ (10)
where $\hat{\theta}$ are the optimized parameters of the original model with training points $\mathcal{X}$ and $\hat{\theta}_{\mathcal{X}^{+},\mathcal{X}^{-}}$ the optimized parameters for training points $\mathcal{X}\cup\mathcal{X}^{+}\setminus\mathcal{X}^{-}$. $\operatorname{Inf}_{f(\boldsymbol{z},\hat{\theta})}$ is the so-called influence function, which is defined by
$\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})$ :-$\coloneq\sum_{i=1}^{m^{+}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{+})-\sum_{i=1}^{m^{-}}\operatorname{Inf}_{f(\boldsymbol{z},\hat{\theta})}(\boldsymbol{x}_{i}^{-}):$ (11)
$\text{with}\quad\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i})$ $\coloneqq\nabla_{\hat{\theta}}f(\boldsymbol{z};\hat{\theta})\cdot H_{\hat{\theta}}^{-1}\cdot\nabla_{\hat{\theta}}L(\boldsymbol{x}_{i};\hat{\theta})\quad\text{.}$ (12)
@@
The optimal parameter under the influence of adding and removing points, is given as

### v2

@@ Abstract
Recently, physics-informed neural networks (PINNs) have emerged as a flexible and promising application of deep learning to partial differential equations in the physical sciences. While offering strong performance and competitive inference speeds on forward and inverse problems, their black-box nature limits interpretability, particularly regarding alignment with expected physical behavior. In the present work, we explore the application of influence functions (IFs) to validate and debug PINNs post-hoc. Specifically, we apply variations of IF-based indicators to gauge the influence of different types of collocation points on the prediction of PINNs applied to a 2D Navier-Stokes fluid flow problem. Our results demonstrate how IFs can be adapted to PINNs to reveal the potential for further studies. The code is publicly available at https://github.com/aleks-krasowski/PINNfluence.
@@ 1 Introduction
Time and time again, deep learning approaches have proven themselves to be exceptionally capable at solving problems that were considered complicated and time-consuming, yielding fast inference times bundled with strong performance. Physics-informed neural networks (PINNs) [28] represent another recent iteration of that trend applied to the realm of partial differential equations (PDEs). By incorporating prior physical knowledge in the form of differential equations [7; 13; 14; 28], they admit a wide range of possible applications such as fluid mechanics, electromagnetics, disease modelling and optics, inter alia [2; 3; 5; 23]. Despite the power of PINNs, there remains the challenge of understanding and improving model behavior when things go wrong [19; 31; 32]. PINNs minimize complex composite losses, which complicates the tracing of poor performance to specific training points or conditions; this is the main motivation for this work.
Influence functions [16] (IFs) systematically assess the contribution of individual training points to a model’s behavior. Recent studies have shown that incorporating physical concepts like temporal causality improves PINN performance [10; 34], highlighting the role of underlying physical principles. By identifying key points, such as those near boundaries or critical regions, we can assess whether the PINN’s learned parameters align with the physics of the problem. For many PDEs, practitioners possess an intuitive understanding of the underlying physics. For example, in magnetostatics, the magnetic field is fundamentally generated by source terms. A well-trained PINN could be expected to reflect this relation. We aim to show how domain experts can use this approach to validate the model by identifying key training points, boundary conditions or physical principles that have the greatest impact on the model’s behavior. This also allows them to estimate the effect of inductive biases and observed data on the model’s predictions. Notably, PINNs also provide a powerful application domain for IFs and other data attribution methods, as models and datasets are usually much smaller than those used in computer vision or natural language processing. This makes the calculation of IFs computationally feasible and the effect of individual data points on the fit of the model is expected to be greater.
@@
  • We adapt influence functions to PINNs as a tool for improving their interpretability.
  • We derive heuristically motivated indicators based on influence functions to showcase how they can be leveraged by domain-knowledgeable users to test and validate PINNs.
  • We empirically demonstrate their usefulness in disambiguating between three different versions of PINNs applied to a Navier-Stokes problem.
@@ 2 Theoretical Background
[28] constitute a machine-learning-based approach to solving PDEs, which are ubiquitous in physical sciences. Let $\Omega\subseteq\mathbb{R}^{n}$ be an open domain and consider the initial boundary value problem (IBVP)
@@
The question being addressed by influence functions [6; 11; 16] is how would the model’s behavior change if certain training points were not present in the training set? By linearly approximating the effect of leave-one-out retraining, IFs provide a possibility of studying a model through the lens of its training data. In order to adapt the influence functions to PINNs we will slightly generalize the approach of [16; 17]. We provide a full derivation and proof, including details on all necessary assumptions, in Section A.1.
Suppose we have trained a model by minimizing a loss $\frac{1}{N}\Sigma_{i}L(\boldsymbol{x}_{i};\theta)$ on a training data set $\mathcal{X}=\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\}$ and derived an optimal parameter $\hat{\theta}\in\Theta$. Furthermore, consider a function $f:A\times\Theta\to\mathbb{R}$, which depends on the parameter $\theta\in\Theta$ and some other parameter $\boldsymbol{z}\in A$ (such as in [16], the loss $L$ for a single test point $\boldsymbol{z}$). Before discussing the case of removing and adding multiple training points, we begin with the case of adding a single training point to $\mathcal{X}$. Thus, we aim to approximate $f(\boldsymbol{z};\hat{\theta})-f(\boldsymbol{z};\hat{\theta}^{+})$ if $\hat{\theta}^{+}$ were optimized over an amended training data set that includes an additional training point $\boldsymbol{x}^{+}$. This can be approximated by
$\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}^{+})\coloneq\nabla_{\theta}f(\boldsymbol{z};\hat{\theta})^{\top}\cdot\mathcal{H}_{\hat{\theta}}^{-1}\cdot\nabla_{\theta}L(\boldsymbol{x}^{+};\hat{\theta})\quad\text{,}$ (3)
where $\mathcal{H}_{\hat{\theta}}\coloneq\nabla^{2}_{\theta}\frac{1}{N}\sum_{i=1}^{N}L(\boldsymbol{x}_{i};\hat{\theta})$ denotes the Hessian w.r.t. $\theta$. This approach can be extended to the addition and removal of multiple training data points: Let $\mathcal{X}^{+}=\{\boldsymbol{x}_{1}^{+},\ldots,\boldsymbol{x}_{m^{+}}^{+}\}$ be the set of training points we want to add and $\mathcal{X}^{-}=\{\boldsymbol{x}_{1}^{-},\ldots,\boldsymbol{x}_{m^{-}}^{-}\}\subseteq\mathcal{X}$ the set we want to remove from our training set. The effect of retraining the model with added and removed training points can then be approximated by ${\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\mathcal{X}^{+},\mathcal{X}^{-})\coloneqq\sum_{i=1}^{m^{+}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{+})-\sum_{i=1}^{m^{-}}\operatorname{Inf}_{f(\boldsymbol{z};\hat{\theta})}(\boldsymbol{x}_{i}^{-})}$. Hence, we can study the influence of single sampling points as well as regions on loss functions, any type of predictions or derived quantities. By studying the influence resulting therefrom, one can compare the model’s behavior with physical intuition. Using this method, we can also estimate the effect of adding and removing entire loss terms for composite losses, as well as several other types of manipulations (see Eqs. 15, 18, 21 and 23 in the appendix).
@@ 3 Experiments
where $\boldsymbol{u}(x,y)\in\mathbb{R}^{2}$ denotes the velocity field, $p(x,y)\in\mathbb{R}$ represents the pressure, $\rho$ is the density, and $\nu$ is the kinematic viscosity. The domain is given by $\Omega=(x_{\min},x_{\max})\times(y_{\min},y_{\max})\setminus C_{r}$, where $C_{r}$ denotes the cylinder with radius $r$. The boundary conditions at the bottom, top, and cylinder are given by $\boldsymbol{u}=0$. The inflow condition on the left has a parabolic profile.
@@
In this study, we aim to answer the following question using influence functions: Can we test whether the PINN has learned certain aspects of the underlying physical principles of the Navier-Stokes problem? Since predictions alone do not reveal the influence of training data or boundary conditions, we devise a set of heuristic indicators.
We propose two IF-based indicators that are designed to capture aspects of the physical processes inherent to the PDE (4) and to serve as exemplary metrics showcasing how expert users can design tests to validate PINNs against their prior knowledge of the problem. As a first indication metric, we concentrate on the flow direction of the Navier-Stokes problem: the inflow is given at the left boundary of the domain and the outflow lies on the opposite side. Figure 2 shows the influence pattern of a training point close to the cylinder.
@@
As a second indicator, we are interested in the fraction of influence that is associated with a given object in the input domain. In the present Navier-Stokes problem, the cylinder plays a central role in the flow field: it obstructs the flow.
@@
Thus, the overall velocity and pressure fields should be heavily influenced by its position. To this end, we devise a region identifier, which is designed to capture the influence associated with a given object or region in the domain (in our case around the cylinder). For the region indicator ($\operatorname{RI}$), which sums up the relevance of removing a region $\Xi\subseteq\overline{\Omega}$, we write:
$\operatorname{RI}(\Xi):=\frac{\sum_{\boldsymbol{x}_{\text{train}}\in\Xi}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}{\sum_{\boldsymbol{x}_{\text{train}}}\left|\operatorname{Inf}_{f(\boldsymbol{x}_{\text{test}};\hat{\theta})}(\boldsymbol{x}_{\text{train}})\right|}\in[0,1]$
A higher value of $\operatorname{RI}$ means that the region in question, such as the cylinder in our case, has an increased influence on the prediction for all other points. This is to be expected for a model that aims to accurately reflect the physics. In this investigation, we choose $\Xi=C_{1.5r}$, meaning that we remove the area corresponding to a circle with 1.5 times the radius of the cylinder, see Fig. 3.
@@ Results
Table 1: Mean values for the Direction $\overline{\operatorname{DI}}$ and Region $\overline{\operatorname{RI}}$ Indicators, aggregated over all $\boldsymbol{x_{\text{test}}}$. A higher score indicates a closer alignment with the indicator’s assumptions.
$\overline{\operatorname{DI}}$ $\overline{\operatorname{RI}}(C_{1.5r})$
$f(x;\theta)$ $\phi_{\text{good}}$ $\phi_{\text{bad}}$ $\phi_{\text{broken}}$ $\phi_{\text{good}}$ $\phi_{\text{bad}}$ $\phi_{\text{broken}}$
@@
We evaluate the three different iterations of PINNs on the Navier-Stokes task. Figures 3(a) to 3(c) show the influence of individual training points on the whole test set, as heatmaps. For $\phi_{\text{good}}$, the most influential points are distributed close to the inflow, on the outflow and especially around the cylinder. This is expected, as these areas determine the fluid flow. Furthermore, we observe that for $\phi_{\text{bad}}$, the upper left boundary is overly influential. In contrast, $\phi_{\text{broken}}$’s influences are again distributed mostly around the cylinder. In Table 1, the aggregated mean values of the indicators are presented. Both $\phi_{\text{good}}$ and $\phi_{\text{bad}}$ show stronger $\operatorname{DI}$ performance. This suggests that both models have successfully learned to capture the dependence of the fluid flow on preceding points. When it comes to the $\operatorname{RI}$ indicator, this trend is even clearer. Here, the good model achieves superior performance, which suggests that the cylinder here has the biggest absolute influence, reflecting its importance for the fluid flow. It is important to note that these indicators are tools for investigating model behavior. Performing well on an indicator does not imply that the model is entirely correct, but rather that it does not fail in the specific aspect the indicator evaluates. Expert knowledge is essential for designing meaningful indicators and further investigation is needed to validate their robustness. Additionally, if the indicator’s assumptions are flawed, the results may be misleading and fail to provide sensible insights.
@@ 4 Conclusion
We showcased how influence functions can help to evaluate PINNs. In particular, we formulated heuristically motivated indicators designed to test certain physical concepts. By applying this method to a Navier-Stokes problem, we observed that the best-performing model consistently aligned with these indicators, confirming their utility in assessing model performance. Still, further exploration of IF-based markers is indicated to rigorously validate the broad applicability of the presented approach. Here, applying novel evaluation frameworks such as [1] could be helpful in future examinations. Using influence functions to improve PINN training through methods like resampling training points [20], enforcing temporal causality [34], or dynamically reweighing loss terms [4] – combined with IF-based training time techniques like TracIN [27] – presents a promising direction for future research. Additionally, further studies on the influence on parts of the loss as well as on different physical quantities, e.g. the vorticity, as well as applying the framework to different PDEs such as the heat equation, Maxwell equations and others appear to be logical next steps.
@@ A.1 Influence Functions
In this part, we aim to formulate the key steps for influence functions, initially introduced in 1974 [11] to study robust estimators, and which were recently applied to neural networks [16; 17]. However, as already observed in [29], the functions typically encountered in neural networks require a more exhaustive analysis. Therefore, we will carefully redraw the results proposed in [16] and generalize them slightly for the application to PINNs. In particular, we will formulate influence functions in a manner that enables the study of the effect of various modifications of the problem (changes in the training set, in the loss weighting or in the PDE) on the prediction and any derived quantity. Furthermore, we will provide rigorous proofs for the statements made in Section 2.
We will start with a crucial lemma for influence functions, which was already noted in [6]. As remarked in [29], influence functions are usually stated for strictly convex functions [^1] Note, that for the application of influence functions with global minima (e.g. [16]), strict convexity is a necessary, but not a sufficient criterion. and relates to global minima. However, in practice, we will typically only reach saddle points [8], which is the reason why the following lemma is extended to stationary points.
Lemma 1.
For $\Theta\subseteq\mathbb{R}^{n}$ and $U\subseteq\mathbb{R}$ open, let $g:\Theta\times U\to\mathbb{R}$ be a twice continuously differentiable function for which there exists an $(\theta_{0},\epsilon_{0})\in\Theta\times U$ such that $\theta_{0}$ is a non-degenerate stationary point of the function $g(\cdot,\epsilon_{0})$, i.e. $\nabla_{\theta}g(\theta_{0},\epsilon_{0})=0$ and the Hessian $\mathcal{H}_{\theta}(\theta_{0},\epsilon_{0})\coloneq\nabla_{\theta}^{2}g(\theta_{0},\epsilon_{0})$ is invertible. Then there exists a non-empty, open set $U_{0}\subseteq U$ with $\epsilon_{0}\in U_{0}$ such that $g(\cdot,\epsilon)$ has a unique stationary point in the neighbourhood of $\theta_{0}$ for each $\epsilon\in U_{0}$ and we call the function describing those stationary points by $h:U_{0}\to\Theta$. This function $h$ is continuously differentiable with derivative
$\frac{\partial h(\epsilon)}{\partial\epsilon}=-\mathcal{H}_{\theta}(h(\epsilon),\epsilon)^{-1}\cdot\left.\nabla_{\theta}\frac{\partial}{\partial\epsilon}g(\theta,\epsilon)\right|_{\theta=h(\epsilon)}\quad\text{.}$ (7)
If $g$ is analytic in a neighbourhood of $(\theta_{0},\epsilon_{0})$, then we can choose $U_{0}$ such that $h$ is analytic as well. Furthermore, let $B\subseteq U$ be an open, connected set with $\epsilon_{0}\in B$, such that for every $\epsilon\in B$ there exists a non-degenerated stationary point of $g(\cdot,\epsilon)$. Then $h$ is extendable to $B$ and this function $h:B\to\Theta$ is continuous differentiable as well with Eq. 7.
@@
By $s:\Theta\times U\to\mathbb{R}^{n}$ we denote the gradient $s(\theta,\epsilon)\coloneq\nabla_{\theta}g(\theta,\epsilon)$, which is continuously differentiable by assumption. Since $(\theta_{0},\epsilon_{0})$ is a stationary point, this will be a root point of $s$. By the implicit function theorem (see e.g. [25, thm. 2]) we can conclude that there is an open set $U_{0}\subseteq U$ around $\epsilon_{0}$ such that there exists a unique function $h:U_{0}\to\Theta$ with $h(\epsilon_{0})=\theta_{0}$ and $s(h(\epsilon),\epsilon)=0$ for all $\epsilon\in U_{0}$. Therefore, for any $\epsilon\in U_{0}$, there is only one unique stationary point in the neighbourhood of $\theta_{0}$. Furthermore, $h$ is continuously differentiable, and its derivative is given by Eq. 7.
If $g$ is analytic, then the same holds for its gradient $s$. By the holomorphic implicit function theorem (see e.g. [9, sec. 7.6]) the function $h$ is also analytic.
If the assumptions of the lemma are satisfied for any $\epsilon\in B$, we can apply the lemma in all of these points. The function $h:B\to\Theta$ is continuous differentiable, since it is continuous differentiable in every point and Eq. 7 extends to all points in $B$. Moreover, $h(B)$ is connected, since $h$ is continuous and $B$ is connected by assumption. ∎
In a general sense, influence functions study the effect of a perturbation of a function, such as a loss function, through a functional, such as the argmin operator. Therefore, we will consider a functional $F:C(\Theta\times U)\times U\to\Theta$ and a continuous function $g:\Theta\times U\to\mathbb{R}^{m}$ with $\Theta\subseteq\mathbb{R}^{n}$ and a neighbourhood $U\subseteq\mathbb{R}$ containing the origin. We assume that $g_{0}(\theta)\coloneq g(\theta,0)$ describes the initial problem, whereas $g(\theta,\epsilon)$ for $\epsilon>0$ corresponds to a perturbed problem, e.g. changes in the training set. The influence function can then be formally expressed as
$\operatorname{Inf}F[g(\cdot,\cdot)]\coloneq\lim_{\epsilon\to 0}\frac{F[g(\cdot,0)]-F[g(\cdot,\epsilon)]}{\epsilon}$ (8)
provided that this limit exists.
In the case of stationary points (in particular for global minima) we obtain a much more convenient expression for the influence function by means of Lemma 1.
Theorem 1.
Let $F$ be the functional of stationary points [^2] To be a functional, $F$ have to map every function to one stationary point or one minimum, respectively. Hence, we might have to adjust e.g. the domain of the input functions to get unique stationary points. (alternatively the argmin operator) and let $g_{0}:\Theta\to\mathbb{R}$ and $\kappa:\Theta\to\mathbb{R}$ be twice continuously differentiable functions. Assume that $g_{0}$ has a non-degenerated, stationary point $\hat{\theta}=F[g_{0}]$ (a global minimum, respectively). Then the influence function is given by
@@
$\operatorname{Inf}F[g_{0}+\epsilon\kappa]=\mathcal{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\kappa(\hat{\theta})\quad\text{,}$ (9)
with the Hessian $\mathcal{H}_{\hat{\theta}}=\nabla_{\theta}^{2}g_{0}(\hat{\theta})$.
Usually we are interested in the effect of a change in the training process on the evaluation of a function $f$, such as the loss or the prediction. Therefore, before proving Theorem 1, we will add the following direct consequence of the definition (8).
Corollary 1.
Let $f:\Theta\to\mathbb{R}^{d}$ be a differentiable function. Then the influence function satisfies a chain rule
$\operatorname{Inf}f(F[g(\cdot,\cdot)])=\nabla_{\theta}f(\theta)\Big{|}_{\theta=F[g(\cdot,0)]}\cdot\operatorname{Inf}F[g(\cdot,\cdot)]\quad\text{.}$ (10)
@@
If the limit in (8) exists, it is nothing else than the derivative. Thus, by applying Lemma 1, the derivative exists and is given by